OpenStudy (anonymous):

The altitude of an equilateral triangle measures 8 √(3). What is the length of one side? _ HELP PLEASE !

OpenStudy (raden):

1/2*s*sqrt(3) = 8sqrt(3) solve for s (side)

OpenStudy (anonymous):

16 square root of 3?

OpenStudy (raden):

nopes... 1/2*s*sqrt(3) = 8sqrt(3) both sides divide by sqrt(3), it can be 1/2*s = 8 please, continue it :)

OpenStudy (anonymous):

lol, thanks (:

OpenStudy (raden):

welcome, the answer is 16, right ?

OpenStudy (anonymous):

yes